OpenStudy (anonymous):

If f(x) = 2x, then f -1(x) = a.-2x b.½x c.x - 2

OpenStudy (anonymous):

ans is b

OpenStudy (anonymous):

what about: If g(x) = 3x - 5, then g -1(x)= a. x+5/3 b. x/3 +5 c. 3x + 5

OpenStudy (anonymous):

ans is a

OpenStudy (anonymous):

do you want explanation or just answers??!! ;)

OpenStudy (anonymous):

sure can you explain

OpenStudy (anonymous):

yes for sure.. here is the method put y=f(x) we get y=3x-5 now solve for x add 5 to both sides we get y+5=3x divide by 3 both sides and then we have x=y+5/3 now remember we had y=f(x) so taking inverse of both sides we get f^-1(y)=x now put value of x we get f^-1(y)=y+5/3 now replace y by x we get f^-1(x)=x+5/3 :)

OpenStudy (anonymous):

thanks a lot can you help with this one? Let R={0, 1, 2, 3} be the range of h(x) = x - 7. The domain of h inverse is a.{-7, -6, -5, -4} b.{7, 8, 9, 10} c.{0, 1, 2, 3}

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

its easy you just have to remember that range of h(x)=domain of h^-1(x) so your ans is c

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

you are most welcome..!!